An investmest of $5600 is made at an annual simple interest of 6.5%. How much additional money must be invested at an annual simple interest rate of 9% so that the total interest earned is 8% of the total investment.
Answer by stanbon(48554)

An investmest of $5600 is made at an annual simple interest of 6.5%. How much additional money must be invested at an annual simple interest rate of 9% so that the total interest earned is 8% of the total investment.
---------------
Amt. invested at 6.5% = $5600 : interest on this is 0.065*5600 = $364.00
Amt invested at 9% = x dollars: interest on this is 0.09x dollars
Amt invested at 8% = (5600+x) ; interest on this is 0.08(5600+x) = 448+0.08x
---------------
EQUATION:
interest + interest = interest
364 + 0.09x = 448+0.08x
0.01x = 84
x = $8400 (Amt that must be invested at 9%)
===============
Cheers,
Stan H.